What is the pH of the best buffer that can be prepared from NH3 and NH4Cl? Kb = 1.8 x 10-5 for NH3.

please show your work
Answer #1

To prepare best buffer, pH should be closest to pKa of the acid.

Given:

Kb = 1.8*10^-5

use:

Ka = Kw/Kb

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Ka = (1.0*10^-14)/Kb

Ka = (1.0*10^-14)/1.8*10^-5

Ka = 5.556*10^-10

Given:

Ka = 5.556*10^-10

use:

pKa = -log Ka

= -log (5.556*10^-10)

= 9.2553

Answer: 9.26